OpenStudy (anonymous):

You move a 75-kg box 35 m. This requires a force of 90 N. How much work is done while moving the box? A. 1,500 N B. 165 N C. 125 N D. 3,150 N

OpenStudy (anonymous):

pwease help? (:

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Hmm so i am gonna say D... ? :D :D lol

OpenStudy (anonymous):

Force = 90N Distance = 35m \[Force \cdot Distance = Work = (90 \cdot 35)J = 3150J\]

OpenStudy (anonymous):

None of the above is correct because they are all listed in Newtons which is not a unit of work.

OpenStudy (anonymous):

Ah, alright. So then Which of the following is required for work to be done on an object? A. In order for work to be done the object must move in the direction of the force and rotate. B. In order for work to be done the object must move over a distance opposite the direction of the force. C. In order for work to be done the object must not move at all. D. In order for work to be done the object must move in the direction of the force and move over a distance. B or C ? (:

OpenStudy (anonymous):

Neither.

OpenStudy (anonymous):

ohhh wait.. but your not thinknig Joules are you? Because

OpenStudy (anonymous):

It said like 55 n -m still equals 55 joules :P If that example makes sense

OpenStudy (anonymous):

In order for work to be done the object must move in the direction of the force for some distance.

OpenStudy (anonymous):

Yes, but you listed the answers as "D. 3,150 N" which is not a \(N\cdot m = J\)

OpenStudy (anonymous):

Hmm right. Well thats odd.. i copied and pasted my question >.<

OpenStudy (anonymous):

Then your worksheet is wrong, or a trick question ;p

OpenStudy (anonymous):

Hmm Prolly a trick question >.< I hate those ;P Thanks though (:

OpenStudy (anonymous):

Probably just the teacher typed it wrong.

OpenStudy (anonymous):

Yea most likely >.<

OpenStudy (anonymous):

Anyway, the answer to your most recent question is D. There must be a force in the direction the object is moving, and (obviously) the object must move.

OpenStudy (anonymous):

Yes, i had gotten that (: You had given me the answer... even though you hadnt said the letter :D Im not too daft to see that ^_^

OpenStudy (anonymous):

Oh, well you hadn't clicked 'good answer' so I wasn't sure if you understood what I meant. ;p

OpenStudy (anonymous):

Haha I need to stop doing that just because the person answers my question... ^_^ No i understood perfectly (: